Find the first partial derivatives of f(x,y,z)=z arctan(y/x) at the point (4, 4, -4).

A. ∂f/∂x(4,4,−4)=
B. ∂f/∂y(4,4,−4)=
C. ∂f/∂z(4,4,−4)=
1
Expert's answer
2020-10-06T18:33:42-0400

A. ∂f/∂x(4,4,−4)= (-4*-4)/(42+42)

=16/32

=1/2

B. ∂f/∂y(4,4,−4)=(-4*4)/(42+42)

=-16/32

=-1/2

C. ∂f/∂z(4,4,−4)=arctan(4/4)

=arctan 1

="\pi" /4